OpenStudy (anonymous):

Express the sum in the simplest form. 3x/x+5 + 7/x+5

OpenStudy (michele_laino):

I s your expression like this: \[\frac{{3x}}{{x + 5}} + \frac{7}{{x + 5}}\]

OpenStudy (anonymous):

@Michele_Laino Yes

OpenStudy (michele_laino):

in that case, since the denominators are equal each other, then you have only to add together 3x and 7, namely: \[\frac{{3x}}{{x + 5}} + \frac{7}{{x + 5}} = \frac{{...}}{{x + 5}}\] please complete

OpenStudy (anonymous):

So then it will just be 3x+7/x+5 ?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thank You

OpenStudy (michele_laino):

:)